each marble bag sold by the company contains 5 blue marbles for every 3 red marbles. if a bag has 24 red marbles how many blue marbles does it contain
Answer by ikleyn(52781)

.
each marble bag sold by the company contains 5 blue marbles for every 3 red marbles.
if a bag has 24 red marbles how many blue marbles does it contain
~~~~~~~~~~~~~


There are  5x  blue marbles and  3x  red marbles,  where x is the common factor.


      3x = 24

       x    = 24/3 = 8.


ANSWER.  5*8 = 40 blue marbles.

Solved and explained.